OpenStudy (anonymous):

3^7*9^5*27^2 the questions is to write the following with only exponent

OpenStudy (valpey):

You mean\(\large\frac{128}{2^3}\), right? What is \(2^3\)?

OpenStudy (anonymous):

the question to write the following with only one exponent

OpenStudy (anonymous):

\[\frac{ 128 }{ 2^{3} }=\frac{ 2^{7} }{ 2^{3} }=2^{7-3}=2^{4}=16\]